# Revisiting Diffusion Q-Learning: From Iterative Denoising to One-Step Action Generation

### v2

@@ 1 Introduction
In recent years, offline reinforcement learning (Offline RL) has achieved impressive progress through the integration of diffusion models, leading to many high-performance algorithms. A prominent example is Diffusion Q-Learning (DQL) (Wang et al., 2022), which replaces the conventional diagonal Gaussian policy in TD3-BC (Fujimoto & Gu, 2021) with a denoising diffusion probabilistic model (DDPM) (Ho et al., 2020). This approach has demonstrated substantial performance gains and has spurred widespread interest in leveraging generative models for policy learning. Notably, DQL remains competitive and often outperforms many more recent methods in both diffusion-based planning and policy optimization (Lu et al., 2025a; Dong et al., 2024).
@@ 2 Related Work
Diffusion Models in Offline Reinforcement Learning. Offline RL aims to learn effective policies from fixed datasets without environment interaction (Levine et al., 2020), with early methods addressing distributional shift via conservative objectives—e.g., CQL (Kumar et al., 2020), TD3+BC (Fujimoto & Gu, 2021), and IQL (Kostrikov et al., 2021). However, these approaches often rely on unimodal Gaussian policies, which struggle to model complex, multi-modal action distributions. To address this, recent approaches have adopted diffusion models (Ho et al., 2020; Song et al., 2020b) for offline RL. These models excel in representing complex distributions and have been applied in various forms: as planners for trajectory generation (Janner et al., 2022; Ajay et al., 2022), as expressive policy networks (Wang et al., 2022; Hansen-Estruch et al., 2023), and as data synthesizers to augment training (Zhu et al., 2023).
@@
Efficient One-Step Diffusion. Our work is inspired by advances in efficient generative modeling with diffusion models and flow-based models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song & Ermon, 2019). To accelerate generation, one line of work focuses on distillation techniques that compress multi-step models into fewer steps (Salimans & Ho, 2022; Sauer et al., 2024; Yin et al., 2024), while another pursues flow matching approaches that learn time-dependent velocity fields for straight-through sampling (Lipman et al., 2022; Liu et al., 2022). Consistency Models (Song et al., 2023) offer another path to one-step generation, but suffer from training instabilities (Song & Dhariwal, 2023; Lu & Song, 2024). More recent methods (Frans et al., 2024; Geng et al., 2025; Zhou et al., 2025) address these limitations by exploiting different physical parametrization using dual time variables, showing improved stability and performance. Leveraging the Mean Flow modeling (Geng et al., 2025), OFQL realizes a one-step flow-based policy while uniquely incorporating the Q-gradient to guide velocity learning, instead of relying solely on supervised learning.
@@ 3.2 Diffusion Q-Learning (DQL)
To form a complete offline RL algorithm, DQL adopts the behavior-regularized actor–critic framework (Wu et al., 2019; Fujimoto & Gu, 2021), alternating between minimizing the actor and critic losses.
@@
where $i\in{1,2}$ indexes the two Q networks for double Q-learning, and $(\phi^{\prime},\theta^{\prime})$ denote target network parameters updated via exponential moving average (EMA) (Fujimoto & Gu, 2021). The actor loss, which focuses on learning the policy, is defined as:
@@ 4 Rationale and methodology
At first glance, diffusion policy sampling appears to be the bottleneck, but resolving DQL’s limitations is far from straightforward. For example, using an efficient solver, e.g., the DDIM solver (Song et al., 2020a), could reduce denoising steps, yet in our experiments, applying DDIM for one-step action generation severely degraded policy performance. Similarly, replacing diffusion with consistency models, such as Consistency-AC (Ding & Jin, 2023), still requires multiple denoising steps and yields lower performance. The most effective one-step approaches are distillation-based methods (Park et al., 2025; Chen et al., 2024; 2023; Lu et al., 2025b), which accelerate inference through student policies but incur an additional training phase or shift the inefficiencies to the distillation stage. This raises a natural question: Can we design a one-step policy that directly eliminates inefficiencies in both training and inference?
@@
Diffusion models generate samples through stochastic and often curved trajectories, which makes one-step sampling challenging. Flow Matching (FM) (Lipman et al., 2022) offers a principled alternative by mapping noise directly to data along smoother, more direct paths (Liu et al., 2022), as illustrated in Fig. 4 (c,d). Modeling the policy via flow matching can potentially improve both efficiency and stability. Let us model the policy as a variant of flow matching based on linear paths and uniform time sampling. Given data $a,s\sim\mathcal{D}$ and noise $\epsilon\sim\mathcal{N}(\mathbf{0},\bm{I})$, FM defines a linear flow path $a_{t}$ and conditional velocity $v_{t}(a_{t}|a;s)$ for a particular $s$ as :
@@
Intuitively, since flows are designed to follow straight trajectories (i.e., ”rectified”) (Liu et al., 2022), modeling the policy via Flow Matching should, in principle, enable one-step generation by setting $\Delta t=1$. However, in practice, the marginal velocity field typically induces a curved overall trajectory, preventing reliable one-step action generation. Importantly, the curvature is not simply a consequence of imperfect neural approximation, but rather an inherent property of the ground-truth marginal velocity field. This phenomenon is illustrated in Fig. 2 (a,b,c).
To enable high-quality one-step generation, we reinterpret the velocity field $v(a_{t},t;s)$ in Flow Matching as the instantaneous velocity, and instead propose to model the average velocity, which directly connects any two arbitrary time steps. Specifically, we define the average velocity over an interval $[r,t]$ as
@@
By approximating $u$ with a neural network, actions can be computed in a single step (i.e., $a=\epsilon-u_{\theta}(\epsilon,r=0,t=1;s)$), thereby eliminating the need for iterative ODE integration in FM. However, computing $u$ from its definition requires integration, which is computationally intractable for optimization. To address this, we adopt an equivalent reformulation based on the MeanFlow Identity (Geng et al., 2025):
@@
where the total derivative expands as $\frac{d}{dt}u(a_{t},r,t;s)=v(a_{t},t;s)\cdot\partial_{a_{t}}u+\partial_{t}u$. Note that the computation of the derivative remains efficient by leveraging Jacobian–vector products (Geng et al., 2025). Notably, when $t=r$, the target $u$ reduces to the instantaneous velocity.
This identity provides a tractable formulation for computing the target average velocity, which can be fitted by a neural network using the following regression loss:
$\mathcal{L}_{FBC}(\theta)=\mathbb{E}_{t,r,r\leq t,(a,s)\sim\mathcal{D},\epsilon}\left\|u_{\theta}(a_{t},r,t;s)-\text{sg}(u_{\text{tgt}})\right\|_{2}^{2}.$ (12)
@@
One-step Flow Q-Learning
To form OFQL, we model the policy with the average velocity parameterization $u_{\theta}$. Its behavior regularization loss is $\mathcal{L}_{FBC}(\theta)$, defined in Eq. 12. In addition, since the target velocity is computed based on the estimate of the instantaneous velocity (11), when sampling $(t,r)$ we enforce a certain flow ratio, i.e., the probability that $t=r$. This design biases training toward learning the instantaneous velocity, while still allowing regression to the average velocity, improving the bootstrapping.
@@ 5 Experiment Setting
Baselines. To rigorously assess OFQL’s performance, we compare it against a broad spectrum of representative baselines, categorized as follows: (1) Non-Diffusion policies: Behavior Cloning (BC), TD3-BC (Fujimoto & Gu, 2021), and IQL (Kostrikov et al., 2021); (2) Diffusion-based planners: Diffuser (Janner et al., 2022), Decision Diffuser (DD) (Ajay et al., 2022); (3) Multi-step Diffusion-based policies: IDQL (Hansen-Estruch et al., 2023), DQL (Wang et al., 2022), and EDP (Kang et al., 2023); and (4) One-step Flow policies: FQL (Park et al., 2025)
@@
HalfCheetah-Medium-Expert 55.2 90.7 86.7 90.3 ± 0.1 88.9 ± 1.9 95.8 ± 0.1 91.3 ± 0.6 96.8 ± 0.3 90.1 ± 2.0 95.2 ± 0.4
@@ 6 Experimental result
Compared to other acceleration approaches, although EDP and IDQL enhance efficiency and stability, they do so at the expense of reducing final performance relative to DQL, whereas OFQL achieves improvements. When compared with one-step FQL (based on distillation), OFQL surpasses it by a significant margin (+14.4). Overall, OFQL offers a superior combination of efficiency and effectiveness, with consistent gains across varying data regimes.
@@
Building on the Q-learning framework, while EDP and IDQL underperform relative to DQL, OFQL achieves a substantial improvement over DQL, raising performance from 64.55 to 84.6 and outperforming all diffusion-based baselines and the one-step FQL. Notably, FQL, which employs a one-step policy for actor–critic training, improves upon DQL from 64.55 to 79.0. We argue that Q-learning is crucial in this domain, and that OFQL and FQL may benefit from avoiding BPTT.
@@ 7 Ablation Study
Table 2 reports the average performance across 9 MuJoCo tasks. DQL+DDIM shows severe degradation (76.3-76.3), suggesting that direct application of improved samplers for one-step inference is ineffective. FBRAC performs better (20.8-20.8), but still lags behind DQL, likely due to the discretization error introduced when performing one-step prediction with curved trajectories. FQL further narrows the gap (9.8-9.8) by distilling a one-step policy from a multi-step flow model. In contrast, OFQL achieves the best results, exceeding DQL by +4.7+4.7 while consistently supporting one-step sampling in both training and inference.
@@
Ablation on flow ratio We study the effect of varying the flow ratio across different datasets in HalfCheetah (Table 5). The best performance is obtained at a flow ratio of $0.5$, achieving $95.2$ on Medium Expert, $63.8$ on Medium, and $52.2$ on Medium Replay. In contrast, using either the flow ratio equal to 1 (equivalent to pure flow matching) or setting it to 0 results in noticeable performance degradation. A moderate flow ratio serves as an effective regularizer, yielding the most stable and robust learning behavior.
@@
Medium Replay 45.2 50.2 52.2 50.3 10.5
Figure 5: D4RL scores across HalfCheetah datasets under varying flow ratios.
@@ Appendix B Architectural and implementation details
Policy Network: The policy is modeled as the average velocity function $u_{\theta}(a_{t},r,t;s)$, where $a_{t}$ denotes the action latent, $t$ and $r$ are timestep variables, and $s$ is the state conditioning input. We adopt the same MLP-based architecture as used in DQL, with the modification of incorporating the additional timestep $r$. Specifically, $u_{\theta}$ is parameterized as a 3-layer multilayer perceptron (MLP) with Mish activations and 256 hidden units per layer. The input to $u_{\theta}$ is the concatenation of the action latent vector, the current state vector, and the sinusoidal positional embeddings of timesteps $t$ and $r$ (time embedding size 64). The output is the predicted average velocity that flows from timestep $t$ to $r$.
Q Networks: We utilize the same Q network architecture as in DQL. Specifically, we employ two Q networks, each implemented as a 3-layer MLP with Mish activations and 256 hidden units per layer. The input to each Q network is the concatenation of the action and the observation, and the output is the estimated state-action value.
@@
Training is conducted for 1000 epochs (2000 for MuJoCo tasks), with each epoch consisting of 1000 gradient steps and a batch size of 256. Both the policy and Q networks are optimized with Adam (Kingma & Ba, 2014), using a learning rate of $3\times 10^{-4}$. For rewards, we adopt the original task rewards in MuJoCo Gym and Kitchen, while following CQL’s modification (Kumar et al., 2020) for AntMaze, consistent with DQL. For evaluation, we report the mean normalized return averaged over three training seeds, with each model evaluated on 150 episodes per task.
Appendix C Flow matching with marginal velocity versus average velocity parameterization on toy datasets
@@
Experiment Setup. As DQL convincingly demonstrates, greater model expressiveness leads to stronger final performance in the actor–critic framework. To illustrate the capability of modeling complex distributions in a one-step setting, we compare Flow Matching with two velocity parameterizations; marginal velocity ($v$-param) and average velocity ($u$-param); across three synthetic datasets: Crescent, Spiral, and Checkerboard. Each dataset challenges the models to capture distinct geometric structures.
@@
Architecture. (1) $v$-param: estimates marginal velocity using a multi-layer perceptron (MLP) with 3 hidden layers (64 units each), with inputs: noise latent, timestep. (2) $u$-param: estimates average velocity using a similar MLP architecture but with an additional target time input. Both models use Mish activations.
Training and Evaluation. We train both models for 100 epochs with a batch size of 2048. The $v$-param is evaluated with varying prediction steps (1, 2, 5, 10), while the $u$-param is evaluated with one denoising step. We visualize the ground truth distribution (blue points) and the generated samples (orange points) on a 2D plot to assess how well each model captures the dataset’s geometric structure.
Results. While the main paper reports results on the Checkerboard dataset, Fig. 7 additionally presents results on two other toy datasets: Spiral and Moon. As illustrated in Figure 7, the $u$-param consistently generates well-structured samples that match the target distribution even in a single step. In contrast, the $v$-param exhibits significant mode collapse and noise in early steps (1–2), requiring up to 10 steps to approximate the target shape. These findings reinforce our main conclusion: modeling the average velocity field enables the $u$-param to achieve accurate and efficient one-step generation, outperforming standard flow models across various geometrically complex distributions.

### v3

# One-Step Flow Q-Learning: Addressing the Diffusion Policy Bottleneck in Offline Reinforcement Learning

@@ 1 Introduction
In recent years, offline reinforcement learning (Offline RL) has achieved impressive progress through the integration of diffusion models, leading to many high-performance algorithms. A prominent example is Diffusion Q-Learning (DQL) (Wang et al., 2022), which replaces the conventional diagonal Gaussian policy in TD3-BC (Fujimoto and Gu, 2021) with a denoising diffusion probabilistic model (DDPM) (Ho et al., 2020). This approach has demonstrated substantial performance gains and has spurred widespread interest in leveraging generative models for policy learning. Notably, DQL remains competitive and often outperforms many more recent methods in both diffusion-based planning and policy optimization (Lu et al., 2025a; Dong et al., 2024).
@@ 2 Related Work
Diffusion Models in Offline Reinforcement Learning. Offline RL aims to learn effective policies from fixed datasets without environment interaction (Levine et al., 2020), with early methods addressing distributional shift via conservative objectives—e.g., CQL (Kumar et al., 2020), TD3+BC (Fujimoto and Gu, 2021), and IQL (Kostrikov et al., 2021). However, these approaches often rely on unimodal Gaussian policies, which struggle to model complex, multi-modal action distributions. To address this, recent approaches have adopted diffusion models (Ho et al., 2020; Song et al., 2020b) for offline RL. These models excel in representing complex distributions and have been applied in various forms: as planners for trajectory generation (Janner et al., 2022; Ajay et al., 2022), as expressive policy networks (Wang et al., 2022; Hansen-Estruch et al., 2023), and as data synthesizers to augment training (Zhu et al., 2023).
@@
Efficient One-Step Diffusion. Our work is inspired by advances in efficient generative modeling with diffusion models and flow-based models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song and Ermon, 2019). To accelerate generation, one line of work focuses on distillation techniques that compress multi-step models into fewer steps (Salimans and Ho, 2022; Sauer et al., 2024; Yin et al., 2024), while another pursues flow matching approaches that learn time-dependent velocity fields for straight-through sampling (Lipman et al., 2022; Liu et al., 2022). Consistency Models (Song et al., 2023) offer another path to one-step generation, but suffer from training instabilities (Song and Dhariwal, 2023; Lu and Song, 2024). More recent methods (Frans et al., 2024; Geng et al., 2025; Zhou et al., 2025) address these limitations by exploiting different physical parametrization using dual time variables, showing improved stability and performance. Leveraging the Mean Flow modeling (Geng et al., 2025), OFQL realizes a one-step flow-based policy while uniquely incorporating the Q-gradient to guide velocity learning, instead of relying solely on supervised learning.
@@ 3.2 Diffusion Q-Learning (DQL)
To form a complete offline RL algorithm, DQL adopts the behavior-regularized actor–critic framework (Wu et al., 2019; Fujimoto and Gu, 2021), alternating between minimizing the actor and critic losses.
@@
where $i\in{1,2}$ indexes the two Q networks for double Q-learning, and $(\phi^{\prime},\theta^{\prime})$ denote target network parameters updated via exponential moving average (EMA) (Fujimoto and Gu, 2021). The actor loss, which focuses on learning the policy, is defined as:
@@ 4 Rationale and methodology
At first glance, diffusion policy sampling appears to be the bottleneck, but resolving DQL’s limitations is far from straightforward. For example, using an efficient solver, e.g., the DDIM solver (Song et al., 2020a), could reduce denoising steps, yet in our experiments, applying DDIM for one-step action generation severely degraded policy performance. Similarly, replacing diffusion with consistency models, such as Consistency-AC (Ding and Jin, 2023), still requires multiple denoising steps and yields lower performance. The most effective one-step approaches are distillation-based methods (Park et al., 2025; Chen et al., 2024; 2023; Lu et al., 2025b), which accelerate inference through student policies but incur an additional training phase or shift the inefficiencies to the distillation stage. This raises a natural question: Can we design a one-step policy that directly eliminates inefficiencies in both training and inference?
@@
Diffusion models generate samples through stochastic and often curved trajectories, which makes one-step sampling challenging. Flow Matching (FM) (Lipman et al., 2022) offers a principled alternative by mapping noise directly to data along smoother, more direct paths (Liu et al., 2022), as illustrated in Figure 4 (c,d). Modeling the policy via flow matching can potentially improve both efficiency and stability. Let us model the policy as a variant of flow matching based on linear paths and uniform time sampling. Given data $a,s\sim\mathcal{D}$ and noise $\epsilon\sim\mathcal{N}(\mathbf{0},\bm{I})$, FM defines a linear flow path $a_{t}$ and conditional velocity $v_{t}(a_{t}|a;s)$ for a particular $s$ as :
@@
Intuitively, since flows are designed so that the overall transport trajectory becomes approximately straight (Liu et al., 2022), one might expect that modeling the policy via Flow Matching could support one-step generation by setting $\Delta t=1$. However, in practice, the sampling trajectory is straight only when the target distribution collapses to a delta distribution or when rectification or similar techniques are explicitly applied. Without such conditions, the marginal velocity field typically induces a curved overall trajectory, preventing reliable one-step action generation. Importantly, the curvature is not simply a consequence of imperfect neural approximation, but rather an inherent property of the ground-truth marginal velocity field. This phenomenon is illustrated in Figure 2 (a,b,c).
To enable high-quality one-step generation, we reinterpret the velocity field $v(a_{t},t;s)$ in Flow Matching as the instantaneous velocity, and instead propose to model the average velocity, which directly connects any two arbitrary time steps. Specifically, we define the average velocity over an interval $[r,t]$ as:
@@
In general, the average velocity is a functional of the instantaneous velocity, i.e., $u=\mathcal{F}[v]$. This field $u$ is fully determined by the instantaneous velocity field $v$ and is independent of any neural network. We therefore treat $u$ as the ground-truth average velocity field and train a neural network $u_{\theta}$ to approximate it using a loss, referred to as the Average-Velocity Matching loss:
$\mathcal{L}_{\text{FBC}^{\star}}(\theta)=\mathbb{E}_{0\leq r\leq t\leq 1;\,s,\epsilon}\!\left[\big\|u_{\theta}(a_{t},r,t;s)-u(a_{t},r,t;s)\big\|^{2}\right].$ (11)
Once $u_{\theta}$ is learned, actions can be generated in a single step through the approximate endpoint map
$a=T_{\theta}(\epsilon,s)=\epsilon-u_{\theta}(\epsilon,r{=}0,t{=}1;s),\qquad\epsilon\sim\mathcal{N}(0,I),$ (12)
which eliminates the iterative ODE integration required by standard Flow Matching. This avoids both the computational overhead and the discretization error associated with numerical ODE solvers. A formal justification for why this one-step procedure preserves the FM action accuracy is provided in Appendix G.
As a result, the learned policy $\pi_{\theta}(a\mid s)$ is the push-forward of the Gaussian prior through the approximate endpoint map:
$\pi_{\theta}=(T_{\theta})_{\#}\mathcal{N}(0,I).$
Moreover, when used as a regularizer, optimizing the average-velocity matching loss $\mathcal{L}_{\text{FBC}^{\star}}(\theta)$ encourages the learned one-step policy $\pi_{\theta}(\cdot\mid s)$ to remain close to the behavior policy $\mu(\cdot\mid s)$, effectively performing behavior cloning. Importantly, this behavior cloning still preserves the ability to model complex, multimodal action distributions through the nonlinear transport map inherited from Flow Matching (see Appendix H for a formal justification).
Practical Loss. In practice, computing $u$ from its definition requires integration, which is computationally intractable for optimization. To address this, we adopt an equivalent reformulation based on the MeanFlow Identity (Geng et al., 2025):
@@
where the total derivative expands as $\frac{d}{dt}u(a_{t},r,t;s)=v(a_{t},t;s)\cdot\partial_{a_{t}}u+\partial_{t}u$. The computation of the derivative also remains efficient by leveraging Jacobian–vector products. Notably, when $t=r$, the target $u$ reduces to the instantaneous velocity.
Equation 13 is mathematically equivalent to Eq. 10 (the detailed derivation is provided in Appendix O). We therefore use the Eq. 13 to compute the average velocity $u_{\text{tgt}}$, avoiding explicit integration in Eq. 10. The resulting field can be fitted by a neural network using the following loss:
$\mathcal{L}_{\text{FBC}}(\theta)=\mathbb{E}_{t,r,r\leq t,(a,s)\sim\mathcal{D},\epsilon}\left\|u_{\theta}(a_{t},r,t;s)-\text{sg}(u_{\text{tgt}})\right\|_{2}^{2}.$ (14)
@@
One-step Flow Q-Learning.
To form OFQL, we model the policy with the average velocity parameterization $u_{\theta}$. Its behavior regularization loss is $\mathcal{L}_{\text{FBC}}(\theta)$, defined in Eq. 14. In addition, since the target average velocity is computed based on the estimate of the instantaneous velocity (Eq. 13), accurate estimation of the instantaneous velocity is crucial for effective average-velocity learning. To encourage this, when sampling $(t,r)$ we enforce a certain flow ratio $\lambda$, i.e., the probability that $t=r$. This design biases training toward learning the instantaneous velocity, while still allowing regression to the average velocity, improving the bootstrapping.
@@ 5 Experiment Setting
Baselines. To rigorously assess OFQL’s performance, we compare it against a broad spectrum of representative baselines, categorized as follows: (1) Non-Diffusion policies: Behavior Cloning (BC), TD3-BC (Fujimoto and Gu, 2021), and IQL (Kostrikov et al., 2021); (2) Diffusion-based planners: Diffuser (Janner et al., 2022), Decision Diffuser (DD) (Ajay et al., 2022); (3) Multi-step Diffusion-based policies: IDQL (Hansen-Estruch et al., 2023), DQL (Wang et al., 2022), and EDP (Kang et al., 2023); and (4) One-step Flow policies: FQL (Park et al., 2025)
@@
HalfCheetah-Medium-Expert 55.2 90.7 86.7 90.3 ± 0.1 88.9 ± 1.9 95.8 ± 0.1 91.3 ± 0.6 96.8 ± 0.3 99.8 ± 0.1 95.2 ± 0.4
@@ 6 Experimental result
Compared to other acceleration approaches, although EDP and IDQL enhance efficiency and stability, they do so at the expense of reducing final performance relative to DQL, whereas OFQL achieves improvements. When compared with one-step FQL (based on distillation), OFQL surpasses it by a significant margin (+13.3). Overall, OFQL offers a superior combination of efficiency and effectiveness, with consistent gains across varying data regimes.
@@
Building on the Q-learning framework, while EDP and IDQL underperform relative to DQL, OFQL achieves a substantial improvement over DQL, raising performance from 64.6 to 84.6 and outperforming all diffusion-based baselines and the one-step FQL. Notably, FQL, which employs a one-step policy for actor–critic training, improves upon DQL from 64.6 to 79.0. We assume Q-learning is crucial in this domain, and that OFQL and FQL may benefit from avoiding BPTT.
@@ 7 Ablation Study
Table 2 reports the average performance across 9 MuJoCo tasks. DQL+DDIM shows severe degradation (76.3-76.3), suggesting that direct application of improved samplers for one-step inference is ineffective. FBRAC performs better (20.8-20.8), but still lags behind DQL, likely due to the discretization error introduced when performing one-step prediction with curved trajectories. FQL further narrows the gap (8.7-8.7) by distilling a one-step policy from a multi-step flow model. In contrast, OFQL achieves the best results, exceeding DQL by (+4.7+4.7) while consistently supporting one-step sampling in both training and inference.
@@
Medium Replay 45.2 50.2 51.2 50.3 10.5
Table 3: D4RL scores across HalfCheetah datasets under varying flow ratios.
Ablation on flow ratio. We study the effect of varying the flow ratio across different datasets in HalfCheetah (Table 3). The best performance is obtained at a flow ratio of $0.5$, achieving $95.2$ on Medium Expert, $63.8$ on Medium, and $52.2$ on Medium Replay. In contrast, using either the flow ratio equal to 1 (equivalent to pure flow matching) or setting it to 0 results in noticeable performance degradation. A moderate flow ratio serves as an effective regularizer, yielding the most stable and robust learning behavior.
@@ 8 Conclusion
9 Acknowledgments
This work was supported by the Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2022-II0951, Development of Uncertainty-Aware Agents Learning by Asking Questions) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-24742969, Intelligent Robotic System using Continual Learning and Multimodal Language Model based Multi Attribute Feedback).
@@ Appendix B Architectural and implementation details
Policy Network: The policy is modeled as the average velocity function $u_{\theta}(a_{t},r,t;s)$, where $a_{t}$ denotes the action latent, $t$ and $r$ are timestep variables, and $s$ is the state conditioning input. We adopt the same MLP-based architecture as used in DQL, with the modification of incorporating the additional timestep $r$. Specifically, $u_{\theta}$ is parameterized as a 3-layer multilayer perceptron (MLP) with Mish activations and 256 hidden units per layer, followed by a linear output layer that maps to the action dimension. The input to $u_{\theta}$ is the concatenation of the action latent vector, the current state vector, and the sinusoidal positional embeddings of timesteps $t$ and $r$ (time embedding size 64). The output is the predicted average velocity that flows from timestep $t$ to $r$.
Q Networks: We utilize the same Q network architecture as in DQL. Specifically, we employ two Q networks, each implemented as a 3-layer MLP with Mish activations and 256 hidden units per layer, followed by a linear output layer that maps to a single action-value dimension. The input to each Q network is the concatenation of the action and the observation, and the output is the estimated state-action value.
@@
Training is conducted for 1000 epochs (2000 for MuJoCo tasks), with each epoch consisting of 1000 gradient steps and a batch size of 256. Both the policy and Q networks are optimized with Adam (Kingma and Ba, 2014), using a learning rate of $3\times 10^{-4}$. For rewards, we adopt the original task rewards in MuJoCo Gym and Kitchen, while following CQL’s modification (Kumar et al., 2020) for AntMaze, consistent with DQL. For evaluation, we follow the protocol of Dong et al. (2024), including the same sampling procedure. We report the mean normalized return averaged over three training seeds, with each model evaluated on 150 episodes per task.
Appendix C DDPM, Flow matching with marginal velocity compared to average velocity parameterization on toy datasets
@@
Figure 7: Comparison of sample quality between DDPM in DQL (left, steps = 1, 2, 5, 10) and u-parameterization in OFQL (right, one step) on toy datasets.
Experiment Setup. As DQL convincingly demonstrates, greater model expressiveness leads to stronger final performance in the actor–critic framework. To illustrate the capability of modeling complex distributions in a one-step setting, we compare (1) Flow Matching with two velocity parameterizations; marginal velocity ($v$-param) and average velocity ($u$-param) (2) DDPM in DQL with average velocity in OFQL; across three synthetic datasets: Crescent, Spiral, and Checkerboard. Each dataset challenges the models to capture distinct geometric structures.
@@
Architecture. (1) $v$-param: estimates marginal velocity using a multi-layer perceptron (MLP) with 3 hidden layers (64 units each), with inputs: noise latent, timestep. (2) $u$-param: estimates average velocity using a similar MLP architecture but with an additional target time input. (3) DDPM: estimate the time step noise using a multi-layer perceptron (MLP) with 3 hidden layers (64 units each), with inputs: noise latent, timestep. All models use Mish activations.
Training and Evaluation. We train both models for 100 epochs with a batch size of 2048 and 40 batches for each epoch. The DDPM, $v$-param is evaluated with varying prediction steps (1, 2, 5, 10), while the $u$-param is evaluated with one denoising step. We visualize the ground truth distribution (blue points) and the generated samples (orange points) on a 2D plot to assess how well each model captures the dataset’s geometric structure.
Results on $v$-param vs $u$-param While the main paper reports results on the Checkerboard dataset, Figure 6 additionally presents results on two other toy datasets: Spiral and Moon. The results show that the $u$-param consistently generates well-structured samples that match the target distribution even in a single step. In contrast, the $v$-param exhibits significant mode collapse and noise in early steps (1–2), requiring up to 10 steps to approximate the target shape. These findings reinforce our main conclusion: modeling the average velocity field enables the $u$-param to achieve accurate and efficient one-step generation, outperforming standard flow models across various geometrically complex distributions.
Results on DDPM vs $u$-param. Figure 7 illustrates that the one-step $u$-parameterization in OFQL achieves expressivity comparable to a full DDPM, despite requiring only a single forward pass. In contrast, DDPM produces highly noisy outputs when restricted to a small number of steps (1–2) and typically requires around 10 denoising steps to approximate the target distribution reliably. Importantly, one-step generation is particularly advantageous in actor–critic RL, as it eliminates the need for backpropagation through time. This makes the one-step $u$-parameterization a more practical choice than DDPM for policy learning, even though both approaches exhibit similar expressive capacity.